# Jason Zhaoxin Fan3, Bo Tang2, Shichao Song1, Mengwei Wang1, Jiawei Yang1

### v1

@@ 1 Introduction
Retrieval-augmented generation (RAG) provides an efficient solution for expanding the knowledge boundaries of large language models (LLMs). Many advanced LLMs, such as ChatGPT OpenAI et al. (2024), Gemini Team et al. (2024), and Perplexy.ai [^1] https://www.perplexity.ai/, have incorporated external retrieval modules within their web platforms. However, during the RAG process, query-relevant texts are processed sequentially through the retriever, the filter before being synthesized into a response by the generator, introducing potential security risks, as attackers can manipulate texts at any stage of the pipeline.
Figure 1: Motivation. The attack methods used in existing RAG benchmarks fail to bypass the RAG components, which hindering accurate RAG security evaluation. Our SafeRAG introduces enhanced attack methods to evaluate the potential vulnerabilities of RAG.
Current attack tasks targeting RAG can be divided into the following four surfaces:
@@
  • Conflict: Knowledge from different sources may conflict with one another, creating opportunities for attackers to manipulate. Simply injecting conflicting texts could prevent LLMs from determining which piece of knowledge is more reliable, resulting in vague or even incorrect responses. Wu et al. (2024a); Liu et al. (2023).
@@
However, most of existing attack tasks at above surfaces often fail to bypass the safety RAG components, making the attacks no longer suitable for RAG security evaluation. There are four main reasons. R-1: Simple safety filters can effectively defend against noise attack Li et al. (2024), as existing noise is often concentrated in superficially relevant contexts, which may actually belong to either similar-topic irrelevant contexts or relevant contexts that do not contain answers. (Fig.1-⑨). R-2: Existing conflict primarily focuses on questions that LLMs can directly answer but contain factual inaccuracies in the related documents Xu et al. (2024). Current adaptive retrievers Tan et al. (2024) have been able to effectively mitigate such context-memory conflict. R-3: Advanced generators demonstrate strong capabilities in detecting and avoiding explicit and implicit toxicity, such as bias, discrimination, metaphor, and sarcasm Sun et al. (2023); Wen et al. (2023). R-4: Traditional DoS attack mainly involves maliciously inserting explicit (Fig.1-⑤⑥) or implicit (Fig.1-⑦⑧) refusal signals into the RAG pipeline. Fortunately, such signals are often filtered out as they inherently do not support answering the question, or they are ignored by generators due to being mixed into evidences Shafran et al. (2024).
@@
Existing benchmarks mainly focus on applying a certain attack task at specific stages of the RAG pipeline and observing the impact of the selected attack on the retriever or generator. In this paper, we introduce the RAG security evaluation benchmark, SafeRAG, which systematically evaluates the potential security risks in the retriever and generator by performing four surfaces of improved attack tasks across different stages of the RAG pipeline. Our main contributions are:
@@
Domain-specific Question,
Comprehensive Question-contexts
@@ 2 Related Work
2.1 RAG Security Benchmark Dataset
Before performing RAG security evaluation, researchers typically design attack datasets meticulously to trigger the vulnerability of RAG. The primary attack types currently include noise, conflict, toxicity, and DoS. As for noise, RGB Chen et al. (2024a) employs a retrieve-filter-classify strategy, dividing the top retrieved contexts related to the query into golden contexts (those containing the correct answer) and relevant noise contexts. RAG Bench Fang et al. (2024) adopts the same approach to construct relevant noise while also introducing irrelevant noise. LRII Wu et al. (2024b) further refines the construction of irrelevant noise through a combination of human effort and LLM assistance, categorizing it into three types: semantically unrelated, partially related, and related to questions.
@@ 2.1 RAG Security Benchmark Dataset
Figure 2: The process of generating attacking texts. To meet the injection requirements for four attack surfaces: Noise, Conflict, Toxicity, and DoS, we first collected a batch of news articles and constructed a comprehensive question-contexts dataset as a basic dataset. Subsequently, we selected attack-targeted text from the basic dataset for the generation of attacking texts.
@@ 2.2 RAG Security Evaluation Metric
In this paper, we focus on both the retrieval safety and generation safety of RAG to provide a more comprehensive evaluation perspective. Moreover, existing RAG safety evaluation methods primarily target open-domain Q&A, with an emphasis on simple fact Q&A in English. We construct a comprehensive question-contexts safety evaluation dataset specifically tailored to the Chinese news domain, aiming to address the complex issues likely to arise in real-world applications.
@@ 3.2 Generation of Comprehensive Question and Golden Contexts
Using DeepSeek [^3] https://www.deepseek.com/, a powerful Chinese LLM engine, and referencing the news title, we generated a comprehensive question and its corresponding 8 pieces of golden contexts for each extracted news segment (Fig.2-②). In total, we obtained 110 unique question-contexts pairs. Additionally, we manually verified and removed data points that did not meet the following criteria: (1) the question is not a comprehensive what, why, or how type question; (2) there are contexts unrelated to the question. Finally, we obtained 100 unique question-contexts pairs, which serve as the basic dataset for attack text generation. The complete generation prompt is detailed in Fig. 12.
3.3 Selection of Attack-Targeted Texts and Generation of Attacking Texts
We select different attack-targeted texts from the question-contexts pairs in the base dataset to generate the specific attacking texts.
@@ 3.3.1 Generation of Silver Noise
To construct silver noise, which includes partial but incomplete answers, we first need to decompose the golden contexts in the base dataset. Specifically, we utilized the knowledge transformation prompt proposed in Chen et al. (2024b) [^4] The complete prompt can be found in Appendix 15. to break the contexts into fine-grained propositions (Fig.2-③), which are the smallest semantic units that are complete and independent as evidence. Then, we selected the proposition with the highest semantic similarity (cosine similarity) to the question as the attack-targeted text, ensuring that the subsequent attack texts achieve a high recall ratio. Finally, we prompted DeepSeek to generate 10 diverse contexts based on the selected attack-targeted text.
@@ 3.3.3 Generation of Soft Ad
For the toxic attack task, we manu ally selected the golden context most susceptible to the injection of malicious soft ads as the attack-targeted text. Then, we encouraged annotators to use the following two methods to create the soft ad attack text by seamlessly incorporating seemingly professional and harmless soft ads into the attack-targeted text: (1) Direct Insertion: Soft ad attack texts are directly inserted into the original context (Fig. 26-①); (2) Indirect Insertion: The original context is modified or adjusted to some extent, for instance, by promoting alongside authoritative entities (such as government, organizations, etc.) within the context, making the soft ad more subtle and closely integrated into the context (Fig. 26-②). The complete annotation requirements are shown in Fig. 17.
@@ 3.4 Attacks on the RAG Pipeline
We perform a final manual verification of the various types of attack contexts obtained and selected appropriate contexts from the golden contexts to construct the final security evaluation datasets. The detailed construction and usage rules for the datasets can be found in Appendix A.1.3.
@@ 4 Evaluation Metrics
4.1 Retrieval Safety Assessment Metric: Retrieval Accuracy
@@
$\text{RA}=\frac{\text{Recall (gc)}+(1-\text{Recall (ac))}}{2},$
@@
4.2 Generation Safety Assessment Metric: F1 Variant and ASR
4.2.1 F1 Variants in Noise and DoS
For silver noise and white DoS tasks, we designed a metric called F1 (avg) to evaluates whether the generated response include fine-grained propositions and thus assess the diversity of the generator. Specifically, we prompt evaluator to complete the multiple-choice question based on the generated response, i.e., the news summary (Fig.7).
@@ 4.2.1 F1 Variants in Noise and DoS
To build multiple options for each data point, we intentionally manipulate certain propositions extracted from the golden contexts to create incorrect ground truth options while retaining the unmodified propositions as correct ground truth options. The evaluator is then tasked with selecting both the correct and incorrect options. Then, we use F1 (correct) to measure the generated response’s ability to correctly identify the correct options, and F1 (incorrect) to evaluate its ability to recognize the incorrect options. We further define F1 (avg), which is the average of F1 (correct) and F1 (incorrect). F1 (avg) quantifies the performance of the generated response in recognizing both correct and incorrect options and reflects the generator’s diversity. If the generator can produce diverse and high-quality responses, it will correctly infer both the correct and incorrect options, resulting in a high F1 (avg).
@@
4.2.2 F1 Variants in Conflict
To construct multiple-choice options for each data point, we deliberately use conflicting facts between contexts as the options and manually annotate the correct ground truth options. As shown in Fig.8, if the response utilizes the correct fact, it can accurately select the correct options. However, selecting the wrong options is not solely due to using conflicting facts; it may also result from generating hallucinated facts. Consequently, the F1 (incorrect) metric is unreliable. Therefore, in the evaluation of inter-context conflict, we only retain the F1 (correct) metric to assess the generator’s performance.
4.2.3 ASR in Conflict, Toxicity, and DoS
@@ 5.1 Settings
The default retrieval window for the silver noise task is set to top K = 6, with a default attack injection ratio of 3/6. For other tasks, the default retrieval window is top K = 2, and the attack injection ratio is fixed at 1/2. We evaluated the impact of using different retrievers (DPR, BM25, Hybrid, Hybrid-Rerank) and filters (OFF, filter NLI Li et al. (2024), compressor SKR Wang et al. (2023)) across different RAG stages (indexing, retrieval, generation) on the contexts retrieved for various generators (DeepSeek, GPT-3.5-turbo, GPT-4, GPT-4o, Qwen 7B, Qwen 14B, Baichuan 13B, ChatGLM 6B). The bold values represent the default settings. Then, we adopt a unified sentence chunking strategy to segment the knowledge base and build the index. The embedding model used is bge-base-zh-v1.5, the reranker is bge-reranker-base, and the evaluator is GPT-3.5-turbo.
@@ 5.2 Results on Noise
We inject different noise ratios into the text accessible in the RAG pipeline, including the knowledge base, retrieved context, and filtered context. As shown in Fig.5, the following observations can be made: (1) Regardless of the stage where noise is injected, the F1 (avg) score exhibits a downward trend as the noise ratio increases, indicating a decline in generation diversity (Fig.5-①); (2) The retriever demonstrates some noise resistance, as noise injected at the knowledge base has approximately 50% chance of not being retrieved. The results in Fig.5-① support this point. Specifically, as the noise ratio increases, the Retrieval Accuracy (RA) of injecting silver noise into the retrieved context or filtered context significantly outperforms that of injecting it into the knowledge base; (3) The performance of injecting noise into the retrieved context and filtered context is similar, indicating that the filter cannot effectively resist silver noise since silver noise still supports answering the query. (4) Different retrievers exhibit varying levels of robustness to noise. Overall, the ranking is Hybrid-Rerank > Hybrid > BM25 > DPR, suggesting that compared to attacking contexts, hybrid retriever and rerankers show a preference for retrieving golden contexts. (5) Compression-based filters like SKR are not sufficiently secure, as they tend to lose detailed information, leading to a decrease in F1 (avg).
@@ 5.3 Results on Conflict, Toxicity, and DoS
(1) After injecting different types of attacks into the texts accessible by the RAG pipeline, it was observed that the retrieval accuracy (RA) and the attack failure rate (AFR) decreased across all three tasks. The ranking of attack effectiveness at different RAG stages was: filtered context > retrieved context > knowledge base. Furthermore, adding conflict attack increased the likelihood of misjudging incorrect options as correct, leading to a drop in F1 (correct). Introducing DoS attack reduced F1 (avg) and severely impacted generative diversity (Fig.6109-①). (2) Retrievers exhibited different vulnerabilities to various attacks. For instance, Hybrid-Rerank was more susceptible to conflict attack, while DPR was more prone to DoS attack. Both experienced a significant decrease in AFR. Additionally, all retrievers showed consistent AFR degradation under toxicity attack. After adding conflict attack, the F1 (correct) scores of all retrievers became similar, indicating stable attack effectiveness. However, DPR was more affected by DoS attack compared to other retrievers, as evidenced by its significantly larger decline in the diversity metric F1 (avg) (Fig.6109-②). (3) The RA of different retrievers was largely consistent across different attack tasks (Fig.6109-③). (4) In conflict tasks, using the SKR filter was less secure because it could compress conflict details, resulting in a decline in F1 (correct). In toxicity and DoS tasks, the NLI filter was generally ineffective, with its AFR close to that of disabling the filter. However, the SKR filter proved to be safe in these tasks, as it was able to compress soft ads and warnings (Fig.6109-④).
@@ 5.4.1 Selection of Generator
We conduct a cumulative analysis of the positive metrics across different attack tasks. As shown in Fig.11, the results show that Baichuan 13B achieved a leading position in multiple attack tasks, particularly excelling in AFR (DoS) and F1 Variants (DoS) metrics [^7] The detailed results are shown in Table 3.. Lighter models are even safer than models such as the GPT series and DeepSeek.

### v2

@@ 1 Introduction
Retrieval-augmented generation (RAG) provides an efficient solution for expanding the knowledge boundaries of large language models (LLMs) Zhao et al. (2024); Gupta et al. (2024); Fan et al. (2024); Wang et al. (2024). Many advanced LLMs, such as ChatGPT OpenAI et al. (2024), Gemini Team et al. (2024), and Perplexy.ai [^1] https://www.perplexity.ai/, have incorporated external retrieval modules within their web platforms. However, in the RAG pipeline, query-relevant texts are processed sequentially through the retriever, the filter before being synthesized into a response by the generator, introducing potential security risks, as attackers can manipulate texts at any stage of the pipeline.
Figure 1: Motivation. The attack tasks considered by the existing RAG benchmarks fail to bypass the RAG components, which hindering accurate RAG security evaluation. Our SafeRAG introduces enhanced attack tasks to evaluate the potential vulnerabilities of RAG.
Current attacks targeting RAG can be divided into four tasks:
@@
  • Conflict: Knowledge from different sources may conflict with one another, creating opportunities for attackers to manipulate. Simply injecting conflict texts could prevent LLMs from determining which piece of knowledge is more reliable, resulting in vague or even incorrect responses. Wu et al. (2024a); Liu et al. (2023); Zou et al. (2024).
@@
However, most of existing attack tasks often fail to bypass the safety RAG components, making the attacks no longer suitable for RAG security evaluation. There are four main reasons. R-1: Simple safety filters can effectively defend against noise attack Li et al. (2024), as existing noise is often concentrated in superficially relevant contexts, which may actually belong to either similar-topic irrelevant contexts or relevant contexts that do not contain answers. (Fig. 1-⑨). R-2: Existing conflict primarily focuses on questions that LLMs can directly answer but contain factual inaccuracies in the related documents Xu et al. (2024). Current adaptive retrievers Tan et al. (2024) have been able to effectively mitigate such context-memory conflict. R-3: Advanced generators demonstrate strong capabilities in detecting and avoiding explicit and implicit toxicity, such as bias, discrimination, metaphor, and sarcasm Sun et al. (2023); Wen et al. (2023). R-4: Traditional DoS attack mainly involves maliciously inserting explicit (Fig. 1-⑤⑥) or implicit (Fig. 1-⑦⑧) refusal signals into the RAG pipeline. Fortunately, such signals are often filtered out as they inherently do not support answering the question, or they are ignored by generators due to being mixed into evidences Shafran et al. (2024).
@@
Existing benchmarks mainly focus on applying a certain attack task at specific stages of the RAG pipeline and observing the impact of the selected attack on the retriever or generator. In this paper, we introduce the RAG security evaluation benchmark, SafeRAG, which systematically evaluates the potential security risks in the retriever and generator by performing four improved attack tasks across different stages of the RAG pipeline. Our main contributions are:
@@
Domain-specific Q&A
News Q&A
@@ 2 Related Work
2.1 RAG Security Evaluation Dataset
Before performing RAG security evaluation, researchers typically design attack datasets meticulously to trigger the vulnerability of RAG (Table 1). The primary attack types currently include noise, conflict, toxicity, and DoS. As for noise, RGB Chen et al. (2024a) employs a retrieve-filter-classify strategy, dividing the top retrieved contexts related to the query into golden contexts (those containing the correct answer) and relevant noise contexts. RAG Bench Fang et al. (2024) adopts the same approach to construct relevant noise while also introducing irrelevant noise. LRII Wu et al. (2024b) further refines the construction of irrelevant noise, categorizing it into types: semantically unrelated, partially related, and related to questions.
@@ 2.1 RAG Security Evaluation Dataset
Figure 2: The process of generating attack texts. To construct SafeRAG dataset covering Noise, Conflict, Toxicity, and DoS, we first collected a batch of news articles and constructed a comprehensive question-contexts dataset as a base dataset. Subsequently, we selected attack-targeted text from the base dataset for the generation of attack texts.
@@ 3.2 Generation of Comprehensive Question and Golden Contexts
Using DeepSeek [^3] https://www.deepseek.com/, a powerful Chinese LLM, and referencing the news title, we generated a comprehensive question and its corresponding 8 pieces of golden contexts for each extracted news segment (Fig. 2-②) [^4] The complete prompt is detailed in Fig. 12.. In total, we obtained 110 unique question-contexts pairs. Then, we manually verified and removed data points that did not meet the following criteria: (1) the question is not a comprehensive what, why, or how type question; (2) there are contexts unrelated to the question. Finally, we obtained 100 unique question-contexts pairs, which serve as the base dataset for attack text generation [^5] More details about base dataset construction can be found in Appendix A.4.2..
3.3 Selection of Attack-Targeted Texts and Generation of Attack Texts
We select different attack-targeted texts from the question-contexts pairs in the base dataset to generate the specific attack texts.
@@ 3.3.1 Generation of Silver Noise
To construct silver noise, which includes partial but incomplete answers, we first need to decompose the golden contexts in the base dataset. Specifically, we utilized the knowledge transformation prompt proposed in Chen et al. (2024b) [^6] The complete prompt can be found in Fig. 15. to break the contexts into fine-grained propositions (Fig. 2-③), which are the smallest semantic units that are complete and independent as evidence. Then, we selected the proposition with the highest semantic similarity (cosine similarity) to the question as the attack-targeted text, ensuring that the subsequent attack texts achieve a high recall ratio. Finally, we prompted DeepSeek to generate 10 diverse contexts based on the selected attack-targeted text [^7] The complete prompt can be found in Fig. 17..
@@ 3.3.3 Generation of Soft Ad
For the toxic attack task, we manu ally selected the golden context most susceptible to the injection of malicious soft ads as the attack-targeted text. Then, we encouraged annotators to use the following two methods to create the soft ad attack text by seamlessly incorporating seemingly professional and harmless soft ads into the attack-targeted text: (1) Direct Insertion: Soft ad attack texts are directly inserted into the original context (Fig. 27-①); (2) Indirect Insertion: The original context is modified or adjusted to some extent, for instance, by promoting alongside authoritative entities (such as government, organizations, etc.) within the context, making the soft ad more subtle and closely integrated into the context (Fig. 27-②) [^9] The annotation requirements are shown in Fig. 18..
@@ 3.4 Attacks on the RAG Pipeline
For each attack task, we integrate attack texts with the golden contexts to construct the SafeRAG dataset [^10] See Appendix A.4.3 for SafeRAG dataset construction.. Using this dataset [^11] The data format of the RAG security evaluation dataset (i.e., SafeRAG dataset) is shown in Fig. 30., we can simulate various attack tasks that RAG may encounter in Q&A tasks. Our threat framework allows attackers to inject attack text at any stage of the RAG pipeline to analyze vulnerabilities under different attacks [^12] See Appendix A.4.4 for threat framework..
@@ 4 Evaluation Metrics
4.1 Retrieval Safety Metric
@@
RA=Recall (gc)+(1Recall ($\text{RA}=\frac{\text{Recall (gc)}+(1-\text{Recall (ac)})}{2},$
@@
4.2 Generation Safety Metric
4.2.1 F1 Variant
Generation security evaluation assesses RAG’s robustness during generation, ensuring accurate and attack-resilient outputs. SafeRAG constructs multiple options for each data point in its dataset, forming a multiple-choice question to test security. During evaluation, the response and the question are fed into the evaluator to obtain results [^13] The prompt for evaluation is provided in Fig. 20..
[…]
Multiple-Choice Construction in Noise and DoS Attacks.
@@ 4.2.1 F1 Variant
In the silver noise and white DoS attack tasks, we construct multiple-choice questions based on fine-grained propositions which are derived by decomposing the golden contexts (Fig. 7). Some of propositions are deliberately distorted by annotators to create incorrect options [^14] The annotation criteria for constructing incorrect options align with those for generating conflicts (Section 3.3.2)., while unmodified propositions serve as correct options.
@@
[…]
Multiple-Choice Construction in Conflict Tasks. In the inter-context conflict task, we have already constructed conflict contexts. Thus, we can simply design multiple-choice questions based on these conflict facts to assess the generator’s decision-making when faced with conflict contexts (Fig. 8). Specifically, we manually label the true or false facts from the conflict contexts as correct and incorrect options, respectively.
[…]
4.2.2 Attack Success Rate (ASR)
@@ 5.1 Settings
The default retrieval window for the silver noise task is set to top K = 6, with a default attack injection ratio of 3/6. For other tasks, the default retrieval window is top K = 2, and the attack injection ratio is fixed at 1/2. We evaluate the security of 14 different types of RAG components against injected attack texts at different RAG stages (indexing, retrieval, and generation), including: (1) retrievers (DPR Reimers and Gurevych (2019), BM25 Robertson and Zaragoza (2009), Hybrid Langchain , Hybrid-Rerank FlagOpen ); (2) filters (OFF, filter NLI Li et al. (2024) [^16] The complete prompt can be found in Fig. 28., compressor SKR Wang et al. (2023) [^17] The complete prompt can be found in Fig. 29.); and (3) generators (DeepSeek, GPT-3.5-turbo, GPT-4, GPT-4o, Qwen 7B, Qwen 14B, Baichuan 13B, ChatGLM 6B). The bold values represent the default settings. Additionally, we adopt a unified sentence chunking strategy to segment the knowledge base during the indexing. The embedding model used is bge-base-zh-v1.5, the reranker is bge-reranker-base.
@@ 5.2 Results on Noise
We inject different noise ratios into the text accessible in the RAG pipeline, including the knowledge base, retrieved context, and filtered context. As shown in Fig. 5, the following observations can be made: (1) Regardless of the stage at which noise injection is performed, the F1 (avg) decreases as the noise ratio increases, indicating a decline in the diversity of generated responses (Fig. 5-①). (2) Different retrievers exhibit varying degrees of noise resistance (Fig. 5-②). The overall ranking of retrievers’ robustness against noise attacks is Hybrid-Rerank > Hybrid > BM25 > DPR. This suggests that hybrid retrievers and rerankers are more inclined to retrieve diverse golden contexts rather than homogeneous attack contexts. (3) As shown in Fig. 5-③, when the noise ratio increases, the retrieval accuracy (RA) for noise injected into the retrieved or filtered context is significantly higher than that for noise injected into the knowledge base. This is because noise injected into the knowledge base has approximately a 50% chance of not being retrieved (Fig. 5-②). (4) The compressor SKR lacks sufficient security. Although it attempts to merge redundant information in silver noise, it severely compresses the detailed information necessary to answer questions within the retrieved context, leading to a decrease in F1 (avg) (Fig. 5-④).
@@ 5.3 Results on Conflict, Toxicity, and DoS
(1) After injecting different types of attacks into the texts accessible at any stage of the RAG pipeline, both F1 (avg) and the attack failure rate (AFR) decline across all three tasks. Specifically, conflict attacks make it difficult for the RAG to determine which information is true, potentially leading to the use of fabricated facts from the attack context, resulting in a drop in metrics. Toxicity attacks cause the RAG to misinterpret disguised authoritative statements as factual, leading to the automatic propagation of soft ads in generated responses, which also contributes to the metric decline. DoS attacks, on the other hand, make the RAG more likely to refuse to answer, even when relevant evidence is retrieved, further reducing the performance metrics. Overall, the ranking of attack effectiveness across different stages is: filtered context > retrieved context > knowledge base (Fig. 6, 9, 10-①). (2) Different retrievers exhibit varying vulnerabilities to different types of attacks. For instance, Hybrid-Rerank is more susceptible to conflict attacks, while DPR is more prone to DoS attacks. The vulnerability levels of retrievers under toxicity attacks are generally consistent (Fig. 6, 9, 10-②). (3) Across different attack tasks, the changes in RA remain largely consistent regardless of the retriever used (Fig. 6, 9, 10-③). (4) In conflict tasks, using the compressor SKR is less secure as it compresses conflict details, leading to a decline in F1 (avg). In toxicity and DoS tasks, the filter NLI is generally ineffective, with its AFR close to that of disabling the filter. However, in toxicity and DoS tasks, the SKR compressor proves to be secure as it effectively compresses soft ads and warning content (Fig. 6, 9, 10-④).
@@ 5.4.1 Selection of Generator
We conduct a cumulative analysis of the positive RAG generation safety metrics across different attack tasks. Fig. 11 shows that Baichuan 13B maintains a leading position in multiple attack tasks, particularly excelling in DoS task [^18] The detailed results are shown in Table 4.. Lighter models are even safer than stronger models such as the GPT series and DeepSeek, as more powerful models may be more sensitive to the toxicity, DoS, and other attacks introduced in this paper.